OpenStudy (anonymous):

what is the solution set for the equation 2x + (the positive value of x) = -2

hartnn (hartnn):

you mean 2+ |x| =-2 ??

hartnn (hartnn):

**2x +|x| = -2

OpenStudy (anonymous):

yes the second way, thanks

hartnn (hartnn):

so, first subtract 2x from both sides, what u get ?

hartnn (hartnn):

correct now if |a| = b then a=b or a=- b so what about -2x-2 = |x| ?

hartnn (hartnn):

confused ? tried ?

OpenStudy (anonymous):

-2?

hartnn (hartnn):

we get 2 equations for -2x-2 = |x| First : -2x-2 = x and 2nd : -2x-2 =-x got this ?

hartnn (hartnn):

so can you solve those individually ?

OpenStudy (anonymous):

-2 and -2/3

hartnn (hartnn):

good work! correct :) now last step is to verify that these are actually the solutions. plug in x=-2 first (and then x=-2/3) in the original equation and verify that left side equals right side or not ?

hartnn (hartnn):

absolutely correct :) so only one solution x=-2 :)

OpenStudy (anonymous):

thanks!

hartnn (hartnn):

welcome ^_^